OpenStudy (anonymous):

If a man walks around a circle 2.125 times, how many metres did he walk if the radius of the circle is 4m?

OpenStudy (anonymous):

help please

OpenStudy (sleepyjess):

Hello! Welcome to OpenStudy! Do you still need help?

563blackghost (563blackghost):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~~~Welcome~to~OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

563blackghost (563blackghost):

Now he walks around a circle meaning the circumfrence now to find the circufrence we would use this equation...\[C=\pi d\] or \[C = \pi r \times 2\]

563blackghost (563blackghost):

Now we know of the radius which is 4 so to find the diameter we would multiply it by 2 so it can equal the entire length of the diameter.... so whats the diameter?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

circumference is 25.12

563blackghost (563blackghost):

Correct! Now we know of the circumfrence! Now he circles it 2.125 times so we simply multiply the total by 2.125.... \[25.12 \times 2.125\]

OpenStudy (anonymous):

53.38 ok

OpenStudy (anonymous):

thank you

563blackghost (563blackghost):

Yup! So the total of meters is 53.38 m^2 xD

563blackghost (563blackghost):

Your Welcome! xD

OpenStudy (anonymous):

i have another question if you have time to help me.

563blackghost (563blackghost):

I can help xD

OpenStudy (anonymous):

tyty

OpenStudy (anonymous):

can i ask here?

563blackghost (563blackghost):

sure :D

OpenStudy (anonymous):

If a ball travels around a circle of radius 4m in 1.5 minutes, what is the angular speed of the ball?

563blackghost (563blackghost):

Hmmm this one is a little hard.... Lets see.... Lets first convert 1.5 minute into seconds.... so there are 60 seconds in one minute and half of a minute is 30 seconds so the total of seconds would be 90..... Now since a whole circle is 2pi radians it would be 2/pi radians .... now I believe we then create a fraction which is setup like this... \[\frac{ 2 \pi }{ 90 }rad/s\] we do this since we want to know what the speed is so we simplify the fraction .... what do you think would be the simplification?

OpenStudy (anonymous):

\[\frac{ \pi }{ 45 }\]

OpenStudy (anonymous):

ty

563blackghost (563blackghost):

Thats what I got as well xD Great Job!

563blackghost (563blackghost):

Im glad I helpful but I must go....byeee xD

OpenStudy (anonymous):

bye thanks a lot